OpenStudy (anonymous):

Solve the systems using substitution. 6x + 2y = 0 4x + y = -1 3x + 2y = -4 x + 2y = -8 5x - 4y = 14 x - 2y = -2 x + 3y = -4 4x + 5y = 5 x + 3y = -5 2x - y = 18

OpenStudy (akashdeepdeb):

Just TAKE X VALUES OF ALL IN TERMS OF Y AND THEN JUST SUBSTITUTE IN THE 2ND EQUATIONS!! :) [TRY]

OpenStudy (anonymous):

these a lot which one you want to solve

OpenStudy (anonymous):

The first one and then i think i could do the rest

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6x+2y=0 -6x= -6x 2y=-6x

OpenStudy (anonymous):

any reply

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

it like you have to leave y by it self

OpenStudy (anonymous):

so the answer is y=-3y

OpenStudy (anonymous):

6x + 2y = 0 subtract 6x on both sides and its 2y=-6x divide each side by 2 so y=-3x plug it in to 4x + y = -1 so 4x + -3x = -1 X=-1 now solve for Y by plugging X in 4(-1) +y=-1 -4+y=-1 Add 4 to both sides y=3

OpenStudy (anonymous):

To test it plug the numbers into 6x + 2y = 0 6(-1) + 2(3) = 0 -6+6=0

OpenStudy (anonymous):

Do you understand how to do the rest? and can I get a medal please :)

OpenStudy (anonymous):

yes